OpenStudy (anonymous):

parallelogram ABCD, angle A is 3x + 12 and angle B is 132°. Show your work in finding the value of x.

hero (hero):

Hint: Adjacent angles of a parallelogram are supplementary.

hero (hero):

Do not give the answer.

hero (hero):

Anyone giving answers will be abuse reported.

OpenStudy (anonymous):

think: 180-132=? ?=3x+12 ?-12=3x now try to solve it

OpenStudy (anonymous):

Really hero.... and 6x?

hero (hero):

2nd Hint: Angle A + Angle B = 180 (3x + 12) + 132 = 180

OpenStudy (anonymous):

that's what i said...

OpenStudy (anonymous):

was it 6? no one told me.....if i was right or wrong...

hero (hero):

@RiyaMoon, use the information I just gave you to solve and use algebra.

OpenStudy (anonymous):

i jsut freaking did and i get six....

hero (hero):

What I mean is, show your steps here.

hero (hero):

Actually type them out

OpenStudy (anonymous):

urg you divide 12 by three and get 2...then the anwser is two.

OpenStudy (anonymous):

nvmforget what i just said. x= 4 so plug 4 in for x and you get 180

hero (hero):

3rd Hint: x does not equal 4

OpenStudy (anonymous):

x=12 then....

hero (hero):

Finally :D

OpenStudy (anonymous):

Exemplary work Hero!

OpenStudy (anonymous):

180-132=48 3x+12=48 subtract 12 on both sides so it would be 3x=36 divide by 3 on both sides getting x=12..........

hero (hero):

Good job, but it would have been better to have written your solution this way: Angle A + Angle B = 180 (3x + 12) + 132 = 180 3x + 12 + 132 = 180 3x + 144 = 180 3x - 144 = 180 3x = 180 - 144 3x = 36 x = 36/3 x = 12 Doing it that way would have shown that you understand that the angles are supplementary and that they have to both add to get a sum of 180.

OpenStudy (anonymous):

so...you just gave the answer...

hero (hero):

Yes, after providing adequate hints and explanations. But more importantly, I was just clarifying a solution the user had already posted. Did you notice that?

OpenStudy (anonymous):

yes...whatever, never mind